ROTATION-TUNNELING ANALYSIS OF EXCITED-STATE PROTON TRANSFER IN DEUTERATED TROPOLONE

\underline{KATHRYN CHEW}, DEACON J. NEMCHICK, JOHN E. WOLFF, and PATRICK H. VACCARO, Department of Chemistry, Yale University, P. O. Box 208107, New Haven, CT 06520-8107 USA.

The origin band of the Ã^{1}B_{2} − X̃^{1}A_{1} (π^{∗} ← π) absorption system in monodeuterated tropolone (TrOD) has been probed with rotational resolution by applying polarization-resolved degenerate four-wave mixing (DFWM) spectroscopy under ambient, bulk-gas conditions. Judicious selection of polarization geometries for incident and detected electromagnetic waves alleviated intrinsic spectral congestion and facilitated dissection of overlapping transitions, thereby enabling refined rotational-tunneling parameters to be extracted for the Ã^{1}B_{2} (π^{∗}π) manifold. A pronounced 2.14(5) cm^{−1} bifurcation of rovibronic features is measured for the zero-point level of electronically excited TrOD, reflecting the presence of a substantial potential barrier along the O−D⋯O ⟷ O⋯D−O reaction coordinate^{a} and representing nearly a ten-fold decrease in magnitude over the analogous tunneling-induced splitting for the parent (TrOH) isotopolog. The dependence of hydron-migration dynamics on internal degrees of freedom will be discussed in light of donor-acceptor displacements incurred by π^{∗} ← π electron promotion and structural effects accompanying selective isotopic modification of the tropolone molecular framework.

aEstimated to be of 1270.6 cm^{−1} height at the fully-optimized EOM-CCSD/aug-cc-pVDZ level of theory; L. A. Burns, D.

Murdock, and P. H. Vaccaro, J .Chem.Phys. 130, 144304 (2009).